OpenStudy (anonymous):

how much will $1,000 amount to after three years if it is invested at 5% interest compounded annually?

OpenStudy (anonymous):

A. $1488.20 B. $1463.27 C. $1352.40 D. $1157.63

OpenStudy (anonymous):

to increase a number by \(5\%\) multiply it by \(1.05\) to do it three times in a row, multiply it by \((1.05)^3\)

OpenStudy (anonymous):

compute with a calculator \[1000(1.05)^3\]

OpenStudy (anonymous):

The equation for that would be B(t)= P(1+r/k)^kt. Meaning balance at time is equal to the initial investment (P) multiplied by 1+ rate divided by amount compounded, to the power of the amount compounded multiplied by time. B(t)=1000(1+0.05/1)^3. It should be $1157.63

OpenStudy (anonymous):

=1000*(1+5%)^3=1157.63 Thus, D should be the right choice.